M is between N and O. Line Segment MO = 15, and line segment MN = 7.6. Find line segment NO.
Answer by ikleyn(52781)

.
M is between N and O. Line Segment MO = 15, and line segment MN = 7.6. Find line segment NO.
~~~~~~~~~~~~~~


From the description, it is obvious that the length of the segment NO is the sum


of the lengths of the segments MO = 15  and MN = 7.6.



ANSWER.  NO = MO + MN = 15 + 7.6 = 22.6 units.